OpenStudy (mcfly47):

Line AB passes through A(-3, 0) and B(-6, 5). What is the equation of the line that passes through the origin and is parallel to line AB ?

OpenStudy (irishboy123):

step 1: get the slope of \(l_{A \to B}\)

OpenStudy (mcfly47):

@IrishBoy123 I worked it out in my notebook and I got -5x – 3y = 0... is that wrong? PS Thanks for the help

OpenStudy (irishboy123):

looks good but why use the negative signs?

OpenStudy (irishboy123):

5x + 3y = 0 is good too

OpenStudy (mcfly47):

mmm because it is parallel?

OpenStudy (mcfly47):

Oh so no negatives then?

OpenStudy (mcfly47):

5x – 3y = 0 -x + 3y = 0 -5x – 3y = 0 3x + 5y = 0 -3x + 5y = 0 these are my options though @IrishBoy123

OpenStudy (irishboy123):

ok weird, but you're right c)

OpenStudy (mcfly47):

Thanks dude!

OpenStudy (irishboy123):

cheers!

OpenStudy (mcfly47):

cheers!